# From non-equilibrium Green functions to Lattice Wigner: A toy model for quantum nanofluidics simulations

### v1

@@ Abstract
Recent experiments of fluid transport in nano-channels have shown evidence of a dramatic reduction of friction due to the coupling between charge-fluctuations in polar fluids and electronic excitations in graphene solids, a phenomenon dubbed ”negative quantum friction”. In this paper, we present a semi-classical mesoscale Boltzmann-Wigner lattice kinetic model of quantum-nanoscale transport and perform a numerical study of the effects of the quantum interactions on the evolution of a one-dimensional nano-fluid subject to a periodic external potential. It is shown that the effects of quantum fluctuations become visible once the quantum length scale (Fermi wavelength) of the quasiparticles becomes comparable to the wavelength of the external potential. Under such conditions, quantum fluctuations are mostly felt on the odd kinetic moments, while the even ones remain nearly unaffected because they are ”protected” by thermal fluctuations. It is hoped that the present Boltzmann-Wigner lattice model and extensions thereof may offer a useful tool for the computer simulation of quantum-nanofluidic transport phenomena.
@@ I Introduction
In the last three decades the Lattice Boltzmann (LB) method has offered a powerful bridge between the atomistic and macroscopic description of flowing matter, with a broad spectrum of applications across many regimes and scales of motion succi2018lattice ; benzi_phys_rep . Although the LB formalism extends to the quantum SUCCI1993327 and relativistic GABBANA20201 realms, its overwhelming body of application is focussed on classical physics, most notably complex fluids and soft matter physrep .
@@
However, the relentless progress of nanotechnology is exposing a growing set of problems whereby quantum phenomena need to be explicitly accounted for, a paradigmatic example being the water flow in carbon nanotubes gao2017nanofluidics . Recently, it has been surmised that quantum interfacial effects in ionic fluids may contribute a sizeable reduction of the drag experienced by water molecules in the proximity of graphene confining walls, a phenomenon called negative quantum friction bocquet1 ; bocquet2 . Such quantum interfacial effects should in principle be treated by ab-initio quantum statistical mechanics methods, such as the non-equilibrium Green function (NEGF) techniques. However, due to its steep computational cost, NEGF is usually replaced by quantum extensions of molecular dynamics wang2009molecular . Even so, reaching to spatial and especially temporal scales of experimental relevance remains a major challenge. There is therefore scope for further coarse-graining, a task at which Lattice Boltzmann has proved very efficient, especially for soft flowing matter applications.
@@
In this paper, we develop a mathematical framework taking from NEGF to LB, and most notably to high-order LB schemes capable of capturing the interplay between classical and quantum non-equilibrium fluctuations, which lies at the heart of quantum nano-fluidic transport, including negative quantum friction effects. More specifically, a high order one-dimensional LB method with third order quantum forcing terms is used to model the evolution of a nano-fluid in the presence of an external periodic potential. Our results suggest that, if the length scale of the quantum force is comparable with that of the external potential, quantum fluctuations are found to disturb odd moments (i.e. current and energy flux) of the distribution functions, whereas such moments are screend from quantum effects if the length scale of the potential is larger. On the contrary, even moments are generally shielded by thermal fluctuations, which prevail over quantum ones.
@@
The paper is organized as follows. In section II we shortly recap the NEGF formalism and its link to the Wigner equation. Sections III and IV are dedicated to discussing the derivation of a high order LB method from the Wigner equation, while section V highlights the application of the method to the realistic case of a hydronic current drive bocquet1 ; bocquet2 . Finally, in section VI we describe the implementation of the one-dimensional LB model with quantum forces and in section VII we present the numerical results, where we study the effect of such forces on the evolution of the power moments of the distribution functions subject to a periodic potential. The main findings and conclusions are summarized in the final section.
II The non-equilibrium Green function
@@
Following Refs. kadanoff ; rammer , we start from a quantum many-body system described by the quantum wavefunction operator $\Psi(x,t)$. The NEGF formalism is based on the Green function associated with the particle generation and destruction operators $\hat{\Psi}$ and $\hat{\Psi}^{+}$
@@
where $1$ and $2$ denote two distinct positions in four-dimensional spacetime $({\bf x},t)$ and subscript -- denotes the anticommutator and brackets denote ensemble averaging over a set of quantum configurations.
By setting ${\bf x}=({\bf x}_{1}+{\bf x}_{2})/2$, ${\bf r}=({\bf x}_{1}-{\bf x}_{2})/2$, $t=(t_{1}+t_{2})/2$ and $s=(t_{2}-t_{1})$ and taking the Fourier-transform, we obtain the associated Wigner function describing the distribution of quasiparticles in eight-dimensional phase-spacetime $({\bf x},{\bf p},t,E)$ wig_physrep
@@
$\Theta=\sum_{k}\left(\frac{\hbar}{2i}\right)^{|k|}\frac{1}{k!}F_{k}\partial_{p}^{k},$ (4)
where $F_{k}=-\partial_{x}^{k}U$ lat_wig , $U$ being the one-body effective potential. The Wigner function bears a close resemblance to a classical probability distribution function, in that its kinetic moments can be associated to the quasiparticle density and current, in close analogy with classical hydrodynamics. This property is key to establish a consistent bridge with the Boltzmann equation. However, its quantum nature is reflected by the fact that $W$ is a pseudo-probability distribution which can take both signs as a result of quantum interference wig_physrep .
@@
quantum interference effects are responsible for the non-positivity of the Wigner function. Eq.(5) is the analogue of the Knudsen number $Kn=\lambda/\delta$, where the molecular mean free path $\lambda$ is replaced by the Fermi wavelength $\lambda_{F}=\frac{\hbar}{mv_{F}}$ (being $\hbar$ the reduced Planck constant and $v_{F}$ the Fermi speed) and $\delta$ is the typical lengthscale. Note that for quadratic potentials, the Wigner function recovers positive-definiteness and becomes fully classical, hence quantum effects are exposed by third order spatial derivatives onward.
@@ III From NEGF to Boltzmann and high-order Lattice Boltzmann
where $Fr$ is the Froude number and $F=-\nabla U$, $U(x,t)$ is the self-interaction potential, i.e. the potential energy acquired by a representative quasiparticle as a result of the interaction with all other quasiparticles. The weak-interaction regime $Fr\ll 1$ permits to associate a single-valued dispersion relation to the quantum excitations, i.e. $\omega=\omega(k)$ and $\gamma=\gamma(k)$, where $k=p/\hbar$ is the wavenumber and $\omega$ and $\gamma$ are the real and imaginary part of the complex wave-frequency. The former controls phase changes (propagation) and the latter amplitude changes (decay/stability). Under such condition the Wigner distribution can be expressed in the so called in-shell representation
@@
$\partial_{t}f+{\bf v}\cdot\nabla_{x}f+{\bf F}\cdot\nabla_{p}f=C(f,f)$ (9)
where $C(f,f)$ is a semiclassical collision operator. In the sequel it proves expedient to replace $C$ with the corresponding single-relaxation time expression bgk
@@
For quantum nanofluidics, there are two additional motivations: first, quantum local equilibria demand energy conservation, hence they need to be formulated on lattices extending beyond the first Brillouin region coelho2014lattice . Second, as discussed earlier on, in the presence of quantum interference, higher order derivatives in momentum space need to be accounted for, which again commends the resort to HOL.
@@ IV Quantum interference and High-Order LB
which gives zero for moments below third order. However, microscopic quantities of order three (i.e. the skewness) couple to the zero-th order moment, which is the fluid density. If, for instance, $\phi(p)=p^{3}/6$, we obtain $\dot{\Phi}_{3}(x)=F_{3}(x)n(x)$; likewise, $\phi(p)=p^{4}/24$ contributes $\dot{\Phi}_{3}(x)=F_{3}(x)J$ and so on. This shows a long-range coupling in momentum space as a result of quantum interference, whence the need of high-order lattices. A detailed list of 2D lattices (whose implementation is challenging but conceptually straightforward) with up to sixteenth order isotropy can be found in sbragaglia2007generalized .
Indeed, previous numerical simulations have shown that the use of HOL leads to more accurate results in the case of anharmonic (fourth-order) potentials, confirming that kinetic moments of order above three do couple to the hydrodynamic sector solorzano2018lattice ; brewer2016lattice . This is because third order derivatives in momentum space, as applied to an Hermite mode of order $n$, excites modes of order $n+3$ in the Hermite ladder.
@@ V Prospective application to hydronic current drive
In this section we discuss the relevant regimes for hydronic current drive nanodevices succi2024keldysh . To convey a concrete idea of a typical application scenario, let us consider a fluid of water molecules flowing in a nano-channel, say a carbon nanotube, confined by carbon walls, either graphite or graphene bocquet1 ; bocquet2 .
@@
Ab-initio analysis based on the quantum non-equilibrium Keldysh formalism predicts that the interaction between water and hydrons in the liquid, electrons and phonons in the solid, leads to a broad variety of energy exchange patterns between the flowing water and the electron-photon ”fluids” in the solid, including the possibility that the electrons may return energy and momentum back to the liquid, thereby leading to a reduction of the friction experienced by water, a mechanism dubbed ”negative quantum friction” bocquet1 ; bocquet2 .
Figure 1: Two different mechanisms driving the electronic current: (a) water molecules subject to charged fluctuations (hydrons, red wavy line) transferring momentum to the electrons (black dots) in the solid through Coulomb interaction; (b) phonons (wavy line), excited by water molecules collisions, drive electrons in the solid. The figure is adapted from Ref.bocquet2 .
@@
By definition, the $q$ and $\mathcal{Q}$ are related via the classical mean free path as $q=\mathcal{Q}{Kn}$. This shows that the condition $\mathcal{Q}>1$ indicates that we are dealing with quantum fluids. To be noted at variance with $q$, which is flow-dependent, that the quantumness is inherently a fluid property. Also to be appreciated that it displays an upper bound dictated by the celebrated AdS-CFT minimum viscosity bound kovtun ; trachenko , which states that any fluid should fulfill the following inequality
@@
Next we consider typical values for a quanto-nanofluidic application, with reference to a nanotube of diameter $D=10$ nm and length $L=100$ nm, with solid wall thickness $a=1$ nm. The electron Fermi wavelength is $\lambda_{F}=h/\sqrt{mE_{f}}\sim 2.5$ nm, where we have taken $m=m_{e}/10$ for the effective electron mass in graphene and $E_{F}\simeq 100$ meV bocquet1 . Assuming longitudinal propagation of the electrons and a transport scale $\delta\sim 10$nm, we have $q\sim 0.25$. This shows that the electron mean free path is comparable with the Fermi wavelength, hence the electronic excitations can be treated semi-classically. We note that in our case, the value of the quantumness $\mathcal{Q}\sim 1$ points indeed to a strongly interacting fluid, but still consistent with the AdS-CFT bound.
@@ VI The D1Q5 model with quantum forces
Here we describe a one-dimensional lattice Boltzmann method to study dynamics of the first five moments of the Wigner distribution function. We consider a D1Q5 lattice consisting of five discrete speeds $c_{i_{x}}=c_{i}$, where $c_{i}=\frac{\Delta x}{\Delta t}$ (with $i=0,1,2,3,4$), $\Delta x$ is the lattice step and $\Delta t$ is the time step, with modulus $c_{0}=0$, $c_{1}=+1$, $c_{2}=-1$, $c_{3}=+2$, $c_{4}=-2$ (see Fig.2).
@@
A set of distribution functions $f_{i}(x,t)$, defined on each site $x$ and time $t$, evolves following a discrete Boltzmann equation succi2018lattice
@@
where $\omega$ is a frequency tuning the relaxation towards the equilibrium and controlling the fluid viscosity $\nu=(1/\omega-1/2)c_{s}^{2}\Delta t$ (with $c_{s}$ lattice sound speed and $c_{s}^{2}=1$), $f_{i}^{eq}$ are the local equilibrium populations and $S_{i}$ are the source terms succi2018lattice . Following common practice in LB theory succi2018lattice ; halim , the former are computed as a second-order Taylor expansion in the fluid velocity $u$ (where $u=u_{x}$) at low Mach number

### v2

# From non-equilibrium Green’s functions to Lattice Wigner: A toy model for quantum nanofluidics simulations

@@ Abstract
Recent experiments of fluid transport in nano-channels have shown evidence of a coupling between charge-fluctuations in polar fluids and electronic excitations in graphene solids, which may lead to a significant reduction of friction a phenomenon dubbed ”negative quantum friction”. In this paper, we present a semi-classical mesoscale Boltzmann-Wigner lattice kinetic model of quantum-nanoscale transport and perform a numerical study of the effects of the quantum interactions on the evolution of a one-dimensional nano-fluid subject to a periodic external potential. It is shown that the effects of quantum fluctuations become visible once the quantum length scale (Fermi wavelength) of the quasiparticles becomes comparable to the lengthscale of the external potential. Under such conditions, quantum fluctuations are mostly felt on the odd kinetic moments, while the even ones remain nearly unaffected because they are ”protected” by thermal fluctuations. It is hoped that the present Boltzmann-Wigner lattice model and extensions thereof may offer a useful tool for the computer simulation of quantum-nanofluidic transport phenomena at scales of engineering relevance.
@@ I Introduction
In the last three decades the Lattice Boltzmann (LB) method has offered a powerful bridge between the atomistic and macroscopic description of flowing matter, with a broad spectrum of applications across many regimes and scales of motion Succi (2018); Benzi et al. (1992). Although the LB formalism extends to the quantum Succi and Benzi (1993) and relativistic Gabbana et al. (2020) realms, its overwhelming body of application is focussed on classical physics, most notably complex fluids and soft matter Tiribocchi et al. (2025).
@@
However, the relentless progress of nanotechnology is exposing a growing set of problems whereby quantum phenomena need to be explicitly accounted for, a paradigmatic example being the water flow in carbon nanotubes Gao et al. (2017); Yu et al. (2023); Lizée et al. (2023, 2024); Bui et al. (2023). Recently, it has been surmised that quantum interfacial effects in dipolar fluids Kavokine et al. (2022); Coquinot et al. (2023) and ionic fluids Coquinot et al. (2024) may contribute a sizeable reduction of the drag experienced by water molecules in the proximity of graphene confining walls, a phenomenon called negative quantum friction which could be crucial for the design of surfaces with low hydrodynamic friction. Such quantum interfacial effects should in principle be treated by ab-initio quantum statistical mechanics methods, such as the non-equilibrium Green’s function (NEGF) techniques. However, due to its steep computational cost, NEGF is usually replaced by quantum extensions of molecular dynamics Wang et al. (2009). Even so, reaching to spatial and especially temporal scales of experimental relevance remains a major challenge. There is therefore scope for further coarse-graining, a task at which LB methods have proved very efficient, especially for soft flowing matter applications.
@@
In this paper, we develop a mathematical framework taking from NEGF to LB, and most notably to high-order LB schemes Chikatamarla and Karlin (2006); Sbragaglia et al. (2007a); Chikatamarla and Karlin (2009); Namburi et al. (2016) capable of capturing the interplay between classical and quantum non-equilibrium fluctuations, which lies at the heart of quantum nano-fluidic transport, including negative quantum friction effects. The main advantage of a numerical approach based on the LB method lies in its capability of reaching space-time scales up two-three times larger than the ones of typical (quantum) molecular dynamics simulations Frenkel and Smit (2001), thus easing the disclosure of new physics emerging at the mesoscale. More specifically, a high order one-dimensional LB method with third order quantum forcing terms is used to model the evolution of a nano-fluid in the presence of an external periodic potential. Our results suggest that, if the length scale of the quantum force is comparable with that of the external potential, quantum fluctuations are found to disturb odd moments (i.e. current and energy flux) of the distribution functions, whereas such moments are screend from quantum effects if the length scale of the potential is larger. On the contrary, even moments are generally shielded by thermal fluctuations, which prevail over quantum ones.
@@
The paper is organized as follows. In section II we shortly recap the NEGF formalism and its link to the Wigner equation. Sections III and IV are dedicated to discussing the derivation of a high order LB method from the Wigner equation, while section V highlights the application of the method to the realistic case of a hydronic current drive Kavokine et al. (2022); Coquinot et al. (2023). Finally, in section VI we describe the implementation of the one-dimensional LB model with quantum forces and in section VII we present the numerical results, where we study the effect of such forces on the evolution of the power moments of the distribution functions subject to a periodic potential. The main findings and conclusions are summarized in the final section.
II The non-equilibrium Green’s function
@@
Following Refs. Kadanoff and Baym (1962); Rammer (2007), we start from a quantum many-body system described by the quantum wavefunction operator $\Psi(x,t)$. The NEGF formalism is based on the Green’s function associated with the particle generation and destruction operators $\hat{\Psi}$ and $\hat{\Psi}^{+}$
@@
where $1$ and $2$ denote two distinct positions in four-dimensional spacetime $({\bf x},t)$ and brackets denote ensemble averaging over a set of quantum configurations.
By setting ${\bf x}=({\bf x}_{1}+{\bf x}_{2})/2$, ${\bf r}={\bf x}_{1}-{\bf x}_{2}$, $t=(t_{1}+t_{2})/2$ and $s=t_{1}-t_{2}$ and taking the Fourier-transform, we obtain the associated Wigner function describing the distribution of quasiparticles in eight-dimensional phase-spacetime $({\bf x},{\bf p},t,E)$ Hillery et al. (1984)
@@
$\Theta=\sum_{k\in N_{odd}}\left(\frac{\hbar}{2i}\right)^{|k|-1}\frac{1}{k!}F_{k}\partial_{p}^{k},$ (4)
where $F_{k}=-\partial_{x}^{k}U$ Solórzano et al. (2018), $U$ being the one-body effective potential. The Wigner function bears a close resemblance to a classical probability distribution function, in that its kinetic moments can be associated to the quasiparticle density and current, in close analogy with classical hydrodynamics. This property is key to establish a consistent bridge with the Boltzmann equation. However, its quantum nature is reflected by the fact that $W$ is a pseudo-probability distribution which can take both signs as a result of quantum interference Hillery et al. (1984).
@@
quantum interference effects are responsible for the non-positivity of the Wigner function. Eq.(5) is the analogue of the Knudsen number $Kn=\lambda/\delta$, where the molecular mean free path $\lambda$ is replaced by the Fermi wavelength $\lambda_{F}=\frac{\hbar}{mv_{F}}$ (being $\hbar$ the reduced Planck constant and $v_{F}$ the Fermi speed) and $\delta$ is the typical lengthscale. Note that for quadratic potentials, the Wigner function recovers positive-definiteness (because the quantum force is identically zero) and becomes fully classical, hence quantum effects are exposed by third order spatial derivatives onward.
@@ III From NEGF to Boltzmann and high-order Lattice Boltzmann
where $Fr$ is the Froude number and $F=-\nabla U$. The weak-interaction regime $Fr\ll 1$ permits to associate a single-valued dispersion relation to the quantum excitations, i.e. $\omega=\omega(k)$ and $\gamma=\gamma(k)$, where $k=p/\hbar$ is the wavenumber and $\omega$ and $\gamma$ are the real and imaginary part of the complex wave-frequency. The former controls phase changes (propagation) and the latter amplitude changes (decay/stability). Under such condition the Wigner distribution can be expressed in the so called in-shell representation
@@
$\partial_{t}f+{\bf v}\cdot\nabla_{x}f+{\bf F}\cdot\nabla_{p}f=C(f,f),$ (9)
where $C(f,f)$ is a semiclassical collision operator. In the sequel it proves expedient to replace $C$ with the corresponding single-relaxation time expression Bhatnagar et al. (1954)
@@
For quantum nanofluidics, there are two additional motivations: first, quantum local equilibria demand energy conservation, hence they need to be formulated on lattices extending beyond the first Brillouin region Coelho et al. (2014). Second, as discussed earlier on, in the presence of quantum interference, higher order derivatives in momentum space need to be accounted for, which again commends the resort to HOL.
@@ IV Quantum interference and High-Order LB
which gives zero for moments below third order. However, microscopic quantities of order three (i.e. the skewness) couple to the zero-th order moment, which is the fluid density. If, for instance, $\phi(p)=p^{3}/6$, we obtain $\dot{\Phi}_{3}(x)=F_{3}(x)n(x)$; likewise, $\phi(p)=p^{4}/24$ contributes $\dot{\Phi}_{3}(x)=F_{3}(x)J$ and so on. This shows a long-range coupling in momentum space as a result of quantum interference, whence the need of high-order lattices. A detailed list of 2D lattices (whose implementation is challenging but conceptually straightforward) with up to sixteenth order isotropy can be found in Sbragaglia et al. (2007b).
Indeed, previous numerical simulations have shown that the use of HOL leads to more accurate results in the case of anharmonic (fourth-order) potentials, confirming that kinetic moments of order above three do couple to the hydrodynamic sector Solorzano et al. (2018); Brewer et al. (2016). This is because third order derivatives in momentum space, as applied to an Hermite mode of order $n$, excites modes of order $n+3$ in the Hermite ladder.
@@ V Prospective application to hydronic current drive
In this section we discuss the relevant regimes for hydronic current drive nanodevices Succi and Montessori (2024). To convey a concrete idea of a typical application scenario, let us consider a fluid of water molecules flowing in a nano-channel, say a carbon nanotube, confined by carbon walls, either graphite or graphene Kavokine et al. (2022); Coquinot et al. (2023).
@@
Ab-initio analysis based on the quantum non-equilibrium Keldysh formalism predicts that the interaction between water and hydrons in the liquid, and electrons and phonons in the solid, leads to a broad variety of energy exchange patterns between the flowing water and the electron-phonon ”fluids” in the solid, including the possibility that the electrons may return energy and momentum back to the liquid, thereby leading to a reduction of the friction experienced by water, a mechanism dubbed ”negative quantum friction” Kavokine et al. (2022); Coquinot et al. (2023).
Figure 1: Two different mechanisms driving the electronic current: (a) water molecules subject to charged fluctuations (hydrons, red wavy line) transferring momentum to the electrons (black dots) in the solid through Coulomb interaction; (b) phonons (wavy line), excited by water molecules collisions, drive electrons in the solid. The figure is adapted from Ref.Coquinot et al. (2023).
@@
By definition, the $q$ and $\mathcal{Q}$ are related via the classical mean free path as $q=\mathcal{Q}{Kn}$. This shows that the condition $\mathcal{Q}>1$ indicates that we are dealing with quantum fluids. To be noted at variance with $q$, which is flow-dependent, that the quantumness is inherently a fluid property. It also displays an upper bound dictated by the celebrated AdS-CFT minimum viscosity bound Kovtun et al. (2005); Trachenko and Brazhkin (2020), which states that any fluid should fulfill the following inequality
@@
Next we consider typical values for a quanto-nanofluidic application, with reference to a nanotube of diameter $D=10$ nm and length $L=100$ nm, with solid wall thickness $a=1$ nm. The electron Fermi wavelength is $\lambda_{F}=h/\sqrt{mE_{f}}\sim 2.5$ nm, where we have taken $m=m_{e}/10$ for the effective electron mass in graphene and $E_{F}\simeq 100$ meV Kavokine et al. (2022). Assuming longitudinal propagation of the electrons and a transport scale $\delta\sim 10$nm, we have $q\sim 0.25$. This shows that the electron mean free path is comparable with the Fermi wavelength, hence the electronic excitations can be treated semi-classically. We note that in our case, the value of the quantumness $\mathcal{Q}\sim 1$ points indeed to a strongly interacting fluid, but still consistent with the AdS-CFT bound.
@@ VI The D1Q5 model with quantum forces
Here we describe a one-dimensional lattice Boltzmann method to study dynamics of the first five moments of the Wigner distribution function. We consider a D1Q5 lattice consisting of five discrete speeds $c_{i_{x}}=c_{i}$, where $c_{i}=\frac{\Delta x}{\Delta t}$ (with $i=0,1,2,3,4$), $\Delta x$ is the lattice step and $\Delta t$ is the time step, and modulus $c_{0}=0$, $c_{1}=+1$, $c_{2}=-1$, $c_{3}=+2$, $c_{4}=-2$ (see Fig.2).
@@
A set of distribution functions $f_{i}(x,t)$, defined on each site $x$ and time $t$, evolves following a discrete Boltzmann equation Succi (2018)
@@
where $\omega$ is a frequency tuning the relaxation towards the equilibrium and controlling the fluid viscosity $\nu=(1/\omega-1/2)c_{s}^{2}\Delta t$ (with $c_{s}$ lattice sound speed and $c_{s}^{2}=1$), $f_{i}^{eq}$ are the local equilibrium populations and $S_{i}$ are the source terms Succi (2018). Following common practice in LB theory Succi (2018); Krüger et al. (2017), the former are computed as a second-order Taylor expansion in the fluid velocity $u$ (with $u=u_{x}$) at low Mach number